OpenStudy (beautifulmystery.):

Estimate the length of the radius, if the area of a circle is 247 mm2. A.9.78 mm B.23.87 mm C.4.43 mm D.8.87 mm

OpenStudy (lfreeman1285):

The equation for the area of a triangle is \[A=\pi r ^{2}\] You were given the area, and pi is a number, so you just have to solve for r

OpenStudy (beautifulmystery.):

circle

OpenStudy (lfreeman1285):

You're looking for the radius, right?

OpenStudy (beautifulmystery.):

of a circle yes not a triangle

OpenStudy (lfreeman1285):

The equation is for the area of a circle.

OpenStudy (lfreeman1285):

If you plug everything in, you'd have \[\sqrt{\frac{ 247 }{ \pi }} = r\] Which simplifies to about 8.87 mm

OpenStudy (beautifulmystery.):

thank you☻

OpenStudy (lfreeman1285):

no problem :)